# ERATTA: Extreme RAG for Table To Answers with Large Language Models

### v2

@@ IV-B Quantitative Response Analysis
The quantitative responses on 500 sample variants of 60 queries generated on the sustainability datasets are shown in Fig. 4. We observe that all responses have average scores around 90%. The number check score is lowest at 89.09% followed by entity check and modifier checks at 94.55% each. Averaged query check and regurgitation check are around 92.73%. This shows the impact of the proposed Extreme-RAG system towards hallucination control and scalability.
@@
IV-C Extreme RAG Extensions
The proposed RAG architecture can further be extended to scenario planning and predictive capabilities using LLMs. Fig. 5 shows a sample flow where a real-time optimization/predictive algorithm is invoked based on the named entities extracted from prompt 1, and the result returned from the algorithm can then be saved into BigQuery followed by a prompt 3 that furnishes the response in natural language. Such extensions/modifications to the code generation LLM prompt (prompt 2) enhance the scalability of the proposed system.
@@ V Conclusion
In this work, we propose a novel multi-LLM system that enables scalable tables-to-answers with an averaged accuracy of $>90\%$. The system flags hallucinations with a series of checks that constantly monitor the multi-LLM outputs. The proposed Extreme-RAG system enables several levels of checks to ensure the appropriate data subset is fetched from large Enterprise-level data tables while returning a natural language response in under 10 seconds per user query. Future works can be directed towards extending RAG approaches to heterogeneous data sources for automation and insight generation tasks.
Acknowledgement
The authors would like to thank Dike Effidua and Bharat Jethwani for all experimentation efforts and Priya Raman for all her leadership.

### v3

@@ IV-B Quantitative Response Analysis
The quantitative responses on 500 sample variants of 60 queries generated on the sustainability datasets are shown in Fig. 4. We observe that all responses have average scores around 90%. The number check score is lowest at 89.09% followed by entity check and modifier checks at 94.55% each. Averaged query-check and regurgitation-check are around 92.73%. This shows the impact of the proposed Extreme-RAG system towards hallucination control and scalability.
@@
IV-C Extreme-RAG Extensions
The extreme-RAG architecture can further be extended to scenario planning and predictive capabilities using LLMs. Fig. 5 shows a sample flow where a real-time optimization/predictive algorithm is invoked based on the named entities extracted from prompt 1, and the result returned from the algorithm can then be saved into BigQuery followed by a prompt 3 that furnishes the response in natural language. Such extensions/modifications to the code generation LLM prompt (prompt 2) enhance the scalability of the proposed system.
@@ IV-C Extreme-RAG Extensions
IV-D Comparison with Agentic-RAG
Langchain (LangGraph) and LlamaIndex have been utilized heavily by several custom products till date to enable LLMs with orchestration capabilities to build intelligent AI agents [11]. For any user-query where the underlying data is relatively large to skim through or highly varying in nature, AI agents can perform a web-search, knowledge identification followed by RAG and automated verification (using Trulens [12]). In Table IV, we compare the performance of an agentic-RAG incorporated using Langchain after prompt 1 to the proposed Extreme-RAG solution. It is noteworthy that agentic-RAG involves prior vector-embedding of the enterprise Tables as an offline process and an additional LLM-based evaluation step.
TABLE IV: Comparison with Agentic RAG after Prompt 1
Method # LLM calls avg time (s) Evaluation metrics Avg score
Proposed 2 8.5 [$s_{1}$-$s_{5}$] 89-95%
Langchain 4-5 40 [Relevance, Toxicity, Grounding] 40-98%
@@ V Conclusion
In this work, we propose a novel multi-LLM system that enables scalable tables-to-answers for Enterprise data tables with an averaged accuracy of $>90\%$ in under 10 seconds per user-query. The proposed process of dividing a user-query for query authentication, query understanding and routing, sub-tablular retrieval followed by answer generation is more stable and faster when compared to agentic-RAG approaches, which makes it a better fit for numeric-intensive use cases such as Sales, Financial and Sustainability domains. Future works can be directed towards extending RAG approaches to heterogeneous data sources for automation and insight generation tasks.